a) If a single card is drawn out of a standard deck of 52 cards, use the appropriate additive counting rule to calculate the probability that it will be a black card or a face card.

b) If two cards are drawn at random, one after another, without replacement, what are the odds in favour that both cards will be a face card?
Answer #1

a)P( probability that it will be a black card or a face card )=P(black card)+P(face card)-P(black and face card) =(26/52)+(12/52)-(6/52) (since there are 26 black, 12 face and 6 black and face card)

=32/52

=8/13

b)

P( both cards will be a face card )=(12/52)*(11/51)=11/221